OpenStudy (anonymous):

SOMEONE PLEASE HELP Solve the proportion x/ x^2 -2= 2/ 3x +4.. What is the value of x? x = −2 x = −2 and x = 2 x = 1 and x = 4 x = 4

OpenStudy (anonymous):

The way you phrase your questions always sounds like you're being molested.

OpenStudy (campbell_st):

\[\frac{x}{x^2 - 2} = \frac{2}{3x + 4}\] so cross mutliply \[x(3x +4) = 2(x^2 - 2)\] you now need to distribute and collect like terms.

OpenStudy (mathstudent55):

Is this the proportion?